Circular No. 5949

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA Sgr 1994 = LUMINOUS RED VARIABLE IN SAGITTARIUS

A. Tomaney and R. M. Rich, Columbia University; R. M. Wagner, Ohio State University; and M. Della Valle, Padua University, report: “Spectra obtained on Mar. 4.54 (IAUC 5944) and 9.54 UT at the Perkins 1.8-m telescope (range 450-570 nm, resolution 0.5 nm) and on Mar. 7.2 at La Silla (IAUC 5945) reveal narrow Balmer emission lines and TiO absorption bands superimposed on a strong red continuum, characteristic of an M giant star. Comparison of the Mar. 4 and 9 spectra shows an evolution from an M0- to an M5-type spectrum associated with more than a magnitude decline in flux. Based on the evolution toward later spectral type and the lack of any emission lines other than the Balmer lines, we propose that this object bears a striking resemblence to a unique superluminous outburst observed in M31 in 1988 (Rich et al. 1989, Ap.J. 341, L51; Mould et al. 1990, Ap.J. 353, L35; Tomaney and Shafter 1992, Ap.J. Suppl. 81, 683). This object brightened by > 5 mag over 2 yr and was discovered at M_{bol} = −10. After the outburst, it faded rapidly, while also changing from M0 to M9; when last detected, the M31 variable was brightest in the infrared L band, and it never showed an obvious nebular-phase spectrum characteristic of a classical nova outburst, suggesting a large ejected mass. Infrared observations (1–10 μm) are urgently required; based on the behavior of the M31 variable, we predict that this new object will rapidly fade in the visible but remain bright in the infrared. The M31 outburst showed dramatic changes in the continuum and emission lines over a few nights. If this outburst is associated with a close binary system, it may exhibit possible pronounced radial-velocity variations associated with the common phase. Higher resolution spectra (resolution 0.1 nm) obtained on Mar. 7.2 and 8.2 at La Silla with the New Technology Telescope (+ EMMI) reveal that Hα exhibits an inverse P-Cyg line profile (correcting the remark on IAUC 5945, line 9). The infall velocity is on the order of 150 km/s (computed from the minimum of the absorption component), and the full width at half-maximum of Hα emission corresponds to ∼ 300 km/s.”

Further photoelectric photometry by A. C. Gilmore (cf. IAUC 5944) at Mt. John Observatory: Mar. 7.71 UT, V = 9.82, B − V = +1.97, V − R = +1.07, V − I = +2.27; 8.70, 10.29, +1.94, +1.20, +2.59; 10.71, 11.18, +1.85, +1.32, +3.09.

1994 March 12 Daniel W. E. Green